OpenStudy (anonymous):

I NEED HELP REALLY BADLY!!! THIS IS THE LAST PART OF MY COURSE! The volume of a cylinder is 72 ft3. The volume of a cube is 36 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube? Provide an explanation for your answer to receive full credit.

OpenStudy (anonymous):

@windsylph @eric_d @RedNeckOutLaw @timo86m @1lorax2 @yomamabf @IdontGetit100 @ore_tee @Preetha @Ashleyisakitty @shyguybros @dsood15 @mathmate @dillonmtorres

OpenStudy (mathmate):

Volume of cylinder = \(\pi r^2 h\) Volume of cone = \(\frac{\pi r^2 h}{3}\) Volume of square based pyramid = \(\large \frac{A_b h}{3}\)

OpenStudy (anonymous):

is that the formula @mathmate ?

OpenStudy (mathmate):

Finally there is Volume of square based prism = \(A_b h\) Yes, these are the formulas needed to do the calculations.

OpenStudy (anonymous):

and how do I do that? @mathmate

OpenStudy (mathmate):

Examine and compare the formulas to find a way to find the answers.

OpenStudy (anonymous):

im horrible at geometry I have no idea how to do this @mathmate

OpenStudy (nikato):

Looking at the formula, The cone is 1/3 the volume of the cylinder so 1/3(72)= ? Then subtract that number from 72 to find the difference

OpenStudy (nikato):

Same with the other. Looking at the formual, the pyramid is 1/3 the volume of the cube, So 1/3(36)=? And that'll be the volume of the square pyramid. Then find the differnce

OpenStudy (anonymous):

I got 48 for the first one then I got 12 @nikato is that the answer?

OpenStudy (nikato):

First one is correct. Remember to substact 12 from 36 for the second. You're suppose to find the difference between the two shapes

OpenStudy (nikato):

Then for the explanation. You can kinda use what I said above And be sure to include the formulas mathmate provided

OpenStudy (anonymous):

ok. so what would I like write for my answer, im not really good at summarizing everything together for my final answer @nikato

OpenStudy (anonymous):

oh lol I sent that right after ok thank you @mathmate and @nikato

OpenStudy (nikato):

I think I pretty much stated it above^ It kinda also includes the explanation